OpenStudy (babyslapmafro):

Please help me finish the following trigonometric substitution (click to see). I have most of the work done just need help towards the end.

OpenStudy (babyslapmafro):

\[\int\limits_{}^{}\sqrt{4-x^2}dx\]

OpenStudy (babyslapmafro):

this is is the result when integrating with respect to theta \[=2\theta+\sin(2\theta)+C\]

OpenStudy (babyslapmafro):

\[x=2\sin(\theta) \] \[\frac{ x }{ 2 }=\sin(\theta)\]

OpenStudy (babyslapmafro):

and here is my triangle....